# Each of those eight coalition logics is also determined by six other kinds of models

### v1

@@ 5.1 Single-coalition-first action models
Definition 15 (Outcome functions and availability functions).
@@
Let $\mathtt{AC}$ be a set of actions and $\mathtt{JA}_{\mathrm{C}}$ be the set of joint actions of $\mathrm{C}$. For any $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}$, if for any $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}=\sigma^{\prime}_{\mathrm{C}}|_{a}$.
@@
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model. The following three conditions are equivalent:
@@
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
@@
      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$, where $\mathrm{C}\neq\emptyset$.
@@
    for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$ and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$:
@@
      $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
    for all $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$:
@@
      $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
@@
      for all fusion $\sigma^{\prime\prime}_{\mathrm{C}}$ of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, for all $t\in\mathtt{ST}$, if $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})$, then $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
@@
Assume (1). Let $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. It suffices to show (2b).
Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\emptyset$. Then, $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{\mathrm{D}}$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. By (1b), $t\in\mathsf{out}_{b}(s,\sigma_{b})$, where $b\in\mathrm{D}$. Then, $t\in\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}$. Note $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{D}=\emptyset$. Similarly, we can show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}\neq\emptyset$ and $\mathrm{D}\neq\emptyset$.
Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\mathrm{D}=\{b_{1},\dots,b_{m}\}$. Let $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$ and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$ and $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$. Then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$.
Note $\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}\cup\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. It suffices to show (1b).
Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{c_{1},\dots,c_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{c_{1}}\cup\dots\cup\sigma_{c_{n}}$.
Note $\mathrm{C}=\{c_{1}\}\cup\mathrm{C}-\{c_{1}\}$. By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ = $\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$.
Note $\mathrm{C}-\{c_{1}\}=\{c_{2}\}\cup\mathrm{C}-\{c_{1},c_{2}\}$. By (2b), $\mathsf{out}_{\mathrm{C}-\{c_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1}\}})$ = $\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2}\}})$.
Finally, we get
$\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap\mathsf{out}_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}}(s,$ $\sigma_{\mathrm{C}}|_{\mathrm{C}-\{c_{1},c_{2},\dots,c_{n-1}\}})$.
That is, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{c_{1}}(s,\sigma_{c_{1}})\cap\mathsf{out}_{c_{2}}(s,\sigma_{c_{2}})\cap\dots\cap$ $\mathsf{out}_{c_{n}}(s,\sigma_{c_{n}})$.
Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
@@
Assume (2). Let $s\in\mathtt{ST}$, $\mathrm{C}\subseteq\mathtt{AG}$, and $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$.
First, we show (3a).
  • Let $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. We want to show $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. We are done if we can show there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. It suffices to show that there is a joint action $\sigma_{\overline{\mathrm{C}}}$ of $\overline{\mathrm{C}}$ such that $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
    Assume $\mathrm{C}=\emptyset$. Then, $\sigma_{\mathrm{C}}=\emptyset$. Then, $w\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Let ${\overline{\mathrm{C}}}=\{a_{1},\dots,a_{n}\}$. Note for every $a\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, there are $\sigma_{a_{1}},\dots,\sigma_{a_{n}}$ such that $w\in\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}}),\dots,w\in\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
    Assume $\mathrm{C}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})$. Then, $w\in\bigcup\{\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\mid\sigma_{a_{1}}\in\mathtt{JA}_{a_{1}}\}$. Then, $w\in\mathtt{suc}(s)$. Assume ${\overline{\mathrm{C}}}=\emptyset$. Then, $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathtt{suc}(s)$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$. Assume ${\overline{\mathrm{C}}}\neq\emptyset$. Assume ${\overline{\mathrm{C}}}=\{b_{1},\dots,b_{n}\}$. As for every $b\in\mathtt{AG}$, $\bigcup\{\mathsf{out}_{b}(s,\sigma_{b})\mid\sigma_{b}\in\mathtt{JA}_{b}\}=\mathtt{suc}(s)$, there are $\sigma_{b_{1}},\dots,\sigma_{b_{n}}$ such that $w\in\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}}),\dots,w\in\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$. Let $\sigma_{\overline{\mathrm{C}}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{n}}$. By (1b), $\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})=\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{n}}(s,\sigma_{b_{n}})$. Then, $w\in\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$.
@@
  • Let $w\in\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$. We want to show $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, there is $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\text{ and}$ $w\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})$. Let $\sigma_{\mathtt{AG}}=\sigma_{\mathrm{C}}\cup\sigma_{\overline{\mathrm{C}}}$. By (2b), $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\overline{\mathrm{C}}}(s,\sigma_{\overline{\mathrm{C}}})$. Then, $w\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$.

### v2

@@ 4.2 Single-coalition-first action models
Definition 15 (Outcome functions and availability functions of single-coalition-first action models).
@@
$\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$
@@
For all $\sigma_{\mathrm{C}},\sigma^{\prime}_{\mathrm{C}},\sigma^{\prime\prime}_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, we say that $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, if for all $a\in\mathrm{C}$, $\sigma^{\prime\prime}_{\mathrm{C}}|_{a}$ equals to $\sigma_{\mathrm{C}}|_{a}$ or $\sigma^{\prime}_{\mathrm{C}}|_{a}$.
The following result offers two alternative definitions of sing-coalition-first action models.
@@
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model and $s\in\mathtt{ST}$.
The following three sets of conditions are equivalent:
@@
      for all nonempty $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a},\text{and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
    It is easy to check this set of conditions is equivalent to the set of constraints on single-coalition-first action models given in Definition 14.
@@
      for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
      for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$;
@@
      for all $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}},\sigma_{\mathrm{C}}^{\prime},\sigma_{\mathrm{C}}^{\prime\prime}\in\mathtt{JA}_{\mathrm{C}}$, if $\sigma^{\prime\prime}_{\mathrm{C}}$ is a fusion of $\sigma_{\mathrm{C}}$ and $\sigma^{\prime}_{\mathrm{C}}$, then $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime}_{\mathrm{C}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma^{\prime\prime}_{\mathrm{C}})$.
@@
Assume (1). We only need to show (2b). Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\sigma_{\mathrm{D}}\in\mathtt{JA}_{\mathrm{D}}$. We want to show $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}=\emptyset$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\emptyset}(s,\emptyset)=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
Assume $\mathrm{C}=\emptyset$ and $\mathrm{D}\neq\emptyset$. Then $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ and $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})=\mathsf{out}_{\emptyset}(s,\emptyset)\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. It suffices to show $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\emptyset}(s,\emptyset)$. Let $t\in\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Let $a\in\mathrm{D}$. By (1b), $t\in\mathsf{out}_{a}(s,\sigma_{\mathrm{D}}|_{a})$. Then, $t\in\bigcup\{\mathsf{out}_{a}(s,\sigma_{a})\mid\sigma_{a}\in\mathtt{JA}_{a}\}=\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $t\in\mathsf{out}_{\emptyset}(s,\emptyset)$. Then, $\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})\subseteq\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$. Then, $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$ $=\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$. Then, $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
The case that $\mathrm{C}\neq\emptyset$ and $\mathrm{D}=\emptyset$ can be handled similarly.
Assume $\mathrm{C}\neq\emptyset$ and $\mathrm{D}\neq\emptyset$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$, $\mathrm{D}=\{b_{1},\dots,b_{m}\}$, $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$, and $\sigma_{\mathrm{D}}=\sigma_{b_{1}}\cup\dots\cup\sigma_{b_{m}}$. By (1b), $\mathsf{out}_{\mathrm{C}\cup\mathrm{D}}(s,\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})\cap\mathsf{out}_{b_{1}}(s,\sigma_{b_{1}})\cap\dots\cap\mathsf{out}_{b_{m}}(s,\sigma_{b_{m}})=\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})\cap\mathsf{out}_{\mathrm{D}}(s,\sigma_{\mathrm{D}})$.
@@
Assume (2). It suffices to show (1b). Let $\mathrm{C}$ be a nonempty subset of $\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. Let $\mathrm{C}=\{a_{1},\dots,a_{n}\}$ and $\sigma_{\mathrm{C}}=\sigma_{a_{1}}\cup\dots\cup\sigma_{a_{n}}$.
Note $\mathrm{C}=\{a_{1}\}\cup\mathrm{C}-\{a_{1}\}$. By (2b), $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\mathsf{out}_{\mathrm{C}-\{a_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1}\}})$.
Note $\mathrm{C}-\{a_{1}\}=\{a_{2}\}\cup\mathrm{C}-\{a_{1},a_{2}\}$. By (2b), $\mathsf{out}_{\mathrm{C}-\{a_{1}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1}\}})=\mathsf{out}_{a_{2}}(s,\sigma_{a_{2}})\cap\mathsf{out}_{\mathrm{C}-\{a_{1},a_{2}\}}(s,\sigma_{\mathrm{C}}|_{\mathrm{C}-\{a_{1},a_{2}\}})$.
Finally, we have $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\mathsf{out}_{a_{1}}(s,\sigma_{a_{1}})\cap\dots\cap\mathsf{out}_{a_{n}}(s,\sigma_{a_{n}})=\bigcap\{\mathsf{out}_{a}(s,\sigma_{{a}})\mid a\in\mathrm{C},\sigma_{{a}}\in\mathtt{JA}_{a}\text{ and }\sigma_{{a}}\subseteq\sigma_{{\mathrm{C}}}\}$.
@@
Assume (2).
@@
First, we show (3a). Let $\mathrm{C}\subseteq\mathtt{AG}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$. We want to show $\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.